# Pair Production in time-dependent Electric field at Finite times

### v4

@@ IV.3 Vacuum Polarisation and depolarisation function
For a better understanding of the phenomenon of particle creation under a strong electric field, we will also trace the evolution of the vacuum polarization function $u(p_{\parallel},t),$ and its counterterm $v(p_{\parallel},t)$, which governs the depolarization. Figure 13 shows the time evolution of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ for different values of $p_{\parallel}$ at zero transverse momentum. Due to pair annihilation being stronger than pair creation in the early times, the depolarisation function $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$[Figs. 13(a) and (c)]. It appears that in the polarization function, there is a sinusoidal function behavior. On the other hand, the depolarization function $v(p_{\parallel},t)$ shows an unimodal Gaussian peak structure in its temporal evolution. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ show oscillations with varying amplitudes during the initial transient stage. These oscillations are particularly pronounced when $p_{\parallel}=0[m]$, as evident from Figs.13(a) and (c). As time progresses, irregular oscillations are observed in the transient stage. However, the oscillations become regular and stable as the system enters the REPP stage. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ now exhibit regular oscillations centered around the zero value. Moreover, as the momentum value $p_{\parallel}$ increases, the amplitudes of these oscillations diminish, as shown in Fig. 13(b) and (d). During the REPP stage, one interesting finding is that $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ demonstrate balancing behavior characterized by similar oscillatory patterns. This balance results from the formation of real electron-positron pairs.
@@ IV.3.1 Momentum Spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$
The longitudinal momentum significantly impacts the vacuum polarization function’s qualitative traits. To understand its dependence on $p_{\parallel}$, we comprehensively analyze the momentum spectra of both $u(p_{\parallel},t)$ and its associated counter term $v(p_{\parallel},t)$. Figure 14 shows the momentum spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ In the initial stages of particle formation, specifically at $t<0$, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ display asymmetric Gaussian peaks with roughly similar profiles. During this time, $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$. Upon closer examination, it becomes evident that the peaks occur at different positions around $p_{\parallel}\approx 1[m]$, as illustrated in Fig. 14(a). At $t=0[m^{-1}]$, the initial Gaussian-like structure of $u(p_{\parallel},t)$ deformed, becoming bi-modal asymmetric structure. The spectrum $u(p_{\parallel},t)$ exhibits a peak at $p_{\parallel}\approx-0.35\ [m]$ and a dip at $p_{\parallel}\approx+0.35\ [m]$. This indicates a significant asymmetry in the function. On the other hand, the $v(p_{\parallel},t)$ spectrum maintains its Gaussian-like unimodal profile, with a peak at $p_{\parallel}=0$. During this stage, while the shape of $u(p_{\parallel},t)$ undergoes considerable changes and becomes asymmetric, the overall form of $v(p_{\parallel},t)$ remains unchanged. Due to the force factor ``$eE(t)"$ and corresponding longitudinal quasi-momentum $P(p_{\parallel},t)=(p_{\parallel}-eA(t))$, the spectrum moved to the left side of the origin, and its peak now located at $p_{\parallel}\approx-2[m]$ for $v(p_{\parallel},t)$ whereas small dip for $u(p_{\parallel},t)$ with some disruption in the tail ( $-1<p_{\parallel}<1$) as shown in the Figs.14(c-d) near to the transient stage.
@@
At $t=24\ [m^{-1}]$, we observe two distinct structures in the polarization and depolarization functions. The first structure appears at $p_{\parallel}=-2\ [m]$ due to earlier particle creation and annihilation events. The second structure, within $-1<p_{\parallel}<1$, exhibits varying amplitude oscillations, with the maximum occurring at zero longitudinal momentum. At this stage, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ compete, as shown in Fig.14(d) and (e). As we approach the REPP stage, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ exhibit a single oscillating structure resembling a sine or cosine function with varying amplitude. The maximum now appears near $p_{\parallel}=0$, dominating compared to the secondary peak (and dip for $u(p_{\parallel},t)$) previously present at $p_{\parallel}\approx-2\ [m]$. As time progresses further, this left-side structure slowly vanishes, as shown in Figs. 14(e) and (f). Eventually, as $A(t)$ reaches a constant value, a balance is achieved between the processes of particle creation and annihilation, as depicted in Fig.14(g). In the late REPP stage, where the particle distribution function $f(p_{\parallel},t)$ is constant, the polarization function $u(p_{\parallel},t)$ is balanced by its counterpart $v(p_{\parallel},t)$. This results in very regular and rapid oscillations with varying amplitude within the Gaussian envelope, as explicitly shown in Figs. 14(h-i). The overall behavior observed in both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ during this process provides valuable insights into the complex dynamics of particle-antiparticle pair creation and annihilation under the influence of the electric field.
IV.3.2 Examining quantum interference effect in LMS of $f(p_{\parallel},t)$: Role of Vacuum Polarization and Depolarization Functions
During the formation of electron-positron pairs from the quantum vacuum, several processes can take place. Electrons are predominantly created and annihilated simultaneously, resulting in electron acceleration and deceleration throughout the formation of real $e^{-}e^{+}$ independent pairs from the virtual $e^{-}e^{+}$ pairs. In these scenarios, the polarization functions $u(p_{\parallel},t)$ and its counterpart $v(p_{\parallel},t)$ are responsible for the acceleration and deceleration of electrons, respectively. In this context, an electron is typically created in the direction of the external field with positive momentum and is subsequently decelerated. As soon as $p_{\parallel}<0$, the electron may be annihilated again. These processes can be understood in terms of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$, which represent the acceleration and deceleration of electrons, respectively.
@@ IV.4 Dependence on Transverse momentum
In this subsection, we extensively investigate the effect of transverse momentum on longitudinal momentum distribution function in dynamical stages of pair production by plotting the time evolution of the LMS for different fixed values of $p_{\perp}$, and the result is displayed in Figure 16. The spectra exhibit a smooth unimodal structure during the initial stages, as depicted in Fig. 16(a). Notably, the peak’s location shifts, as illustrated in Fig.16(b), accordingly with the $(p_{\parallel}-eA(t))=0$ as explained in section IV.2. Within a narrow spectral region for $-1<p_{\parallel}<1,$ some irregular oscillating structure is observed. Furthermore, these figures indicate that the peak value of the spectra decreases with the increase in the value of transverse momentum. Fig. 16 (c), at $t=30[m^{-1}]$ spectra shows bimodal distribution. The central peak structure with onset oscillation located at $p_{\parallel}=0$ dominates over another peak at $p_{\parallel}=-2[m]$ for a small value of $p_{\perp}.$ One interesting fact observed that the central peak magnitude diminished very rapidly compared to other peaks and that oscillating structure does not exist for high transverse momentum at this time. The existence of oscillation in the central peak depends on what time the particle reaches the on-shell configuration from the initial off-shell mass configuration, which critically depends on both transverse as well as longitudinal momentum values explained in section IV.1.
@@
Overall, we know that from  section IV.1 increasing the momentum value particle takes a longer time to reach in REPP stage. So, depending upon the value of $(p_{\parallel},p_{\perp})$, we can say that different portions of the spectra point to different dynamical stages. The insets of Fig. 16 (c) shows only an irregular oscillatory structure in the tail region of the spectra for $p_{\perp}=0.8[m]$, and the central peak structure is not observed. This absence is attributed to the fact that particles are in the QEPP stage for higher transverse values. At $t=50[m^{-1}],$ for $p_{\perp}=0$, that oscillation is nearly smooth, but for high transverse value,$p_{\perp}=0.8[m]$ magnitude of peak diminished see Fig.16 (d), however as shown in insert figure that onset oscillation still observed only magnitude get decreased. From this observation, we can say that the central peak structure does not show the same behavior for a higher value of $p_{\perp}$. This is exactly what is implied by Fig. 16 (d). As we see that for higher transverse momentum, we still see that central peak with onset oscillation, which means it just reaches the REPP region. We also observe that for a higher value of $p_{\perp}$ width of the spectra, $\Delta p_{\parallel}$ gets decreased, see insert Fig. 16 (d).
Although a detailed analysis is beyond the scope of the present manuscript, the time scales are modified in the presence of transverse momentum $p_{\perp}$ as follows ( the data not shown):
$t_{cp}$ $=0.4606\tau^{1.141}{E_{0}}^{-1.273}{p_{\perp}}^{0.648},$ (104)
$t_{sep}$ $=0.7808\tau^{1.125}{E_{0}}^{-1.048}{p_{\perp}}^{0.536841},$ (105)
$t_{dis}$ $=2.929\tau^{0.85365}{E_{0}}^{-0.830}{p_{\perp}}^{0.3182}.$ (106)
These equations highlight how the time scales are influenced by the electric field strength $E_{0}$, the pulse duration $\tau$, and the transverse momentum $p_{\perp}$. The time scales found in Ref. Diez et al. (2023) differ from the scaling we have identified. In our analysis, we demonstrate that these time scales also depend on the transverse momentum $p_{\perp}$, a dependence that was absent in Ref. Diez et al. (2023).
@@ IV.5 Transverse Momentum Spectrum
Following our exploration of the impact of transverse momentum on the longitudinal momentum distribution function, we now shift our attention to the transverse momentum distribution function. Figure 17 shows the time evolution of the TMS for created particles at different times. At the initial time $t=-5\,[m^{-1}]$, TMS exhibits a smooth, unimodal structure with a peak at $p_{\perp}=0[m]$, where the electric field is at $75\%$ of its maximum strength. As the electric field reaches its maximum value at $t=0\,[m^{-1}]$, the spectrum undergoes a dramatic change: both the peak height and width of the spectrum significantly shift, although the peak remains at $p_{\perp}=0[m]$. After the peak reaches its maximum height, it rapidly decreases as the electric field weakens.
The single peak structure initially observed at $p_{\perp}=0$ has disappeared, and new peaks have emerged at $p_{\perp}\approx\pm 0.46\,[m]$ in TMS. As an example, at $t=7\,[m^{-1}]$, TMS exhibits a bimodal profile, as shown in Fig. 17(c). As time progresses, the peaks at $p_{\perp}\approx\pm 0.46\,[m]$ tend to merge, leading to the formation of a peak at $p_{\perp}=0[m]$ with two weakly pronounced peaks, as illustrated in Fig. 17(d). Those weakly pronounced peaks grow over time, while the height of the central peak decreases. As a result, TMS again exhibits a multi-modal structure, and $f(p_{\perp},t)$ reaches its maximum value at non-zero transverse momentum, specifically at $p_{\perp}\approx\pm 0.65\,[m]$ at $t=12\,[m^{-1}]$, with a small peak still present at $p_{\perp}=0[m]$, as shown in Fig. 17(e). Overall, TMS displays a substructure profile where the dominant peak is governed by non-zero values of $p_{\perp}$. Upon closer inspection, TMS reveals that the peak at $p_{\perp}=0$ becomes invisible again, and the locations of the dominant side peaks undergo slight modulation at $t=14\,[m^{-1}]$. Specifically, the dominant peak at $p_{\perp}=0.67[m]$ splits into two weakly pronounced peaks: one at $p_{\perp}=0.45[m]$ and the other at $p_{\perp}=0.94[m]$, as depicted in Fig. 17(f). As time progresses, the maximum peak at zero transverse momentum reappears. Although this central peak stands out, there are also side peaks (not shown in the figure). These side peaks eventually dominate as the central peak fades, just like in the earlier stages. In this way, the side peaks and the central peak reach the same height, as shown in the figure at $t=16.6[m^{-1}]$, and the spectrum shows a three-peak pattern.
We can say that the spectrum of created particles at a given time exhibits key features, indicating that the spectrum undergoes continuous changes and reveals two distinct structures. One scenario shows a maximum peak at $p_{\perp}=0[m]$, suggesting particles with no transverse momentum. The other case displays a maximum peak of the spectrum at a non-zero $p_{\perp}$ value. In both cases, there are multiple peaks, showing the complex behavior of the spectrum over time, as seen in figure 17(a)-(h).
@@
Figure 17: TMS of created particles in the presence of time-dependent Sauter pulse at different times. The longitudinal momentum is considered to be zero, and all the units are taken in the electron mass units.The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$
@@
The effect of created pairs tends to have a peak in the momentum distribution function at non-zero transverse momentum, as discussed by Krajewska et al. Krajewska and Kamiński (2019). It is demonstrated that the transverse momentum distribution of real pairs exhibits an off-axis maximum above the one-photon threshold frequency (or high frequency); otherwise, the maximum for the transverse distribution function occurs at $p_{\perp}=0[m]$ in the presence of an oscillating time-dependent electric field. This study gives the clue that our case of off-axis maxima or multi-modal structure may also be observed due to multiphoton transition processes. To confirm this, we analyze the probability of pair production by calculating the amplitude in the two-level equivalent formalism Krajewska and Kamiński (2019). We decompose the time-dependent electric field $E(t)$ via the Fourier transformation into cosine and sine components. This decomposition isolates individual frequency components of the field, allowing us to study their effects on pair production at finite times. For a given frequency $\nu$, the corresponding electric field amplitude $\tilde{E}(\nu)$ facilitates the evaluation of the particle production spectrum at a specific finite time $t$. The spectrum exhibits multiple peaks, forming a multimodal profile. Each peak corresponds to distinct energy or momentum exchanges between the field and the particle pairs, characteristic of multi-photon absorption. By considering different frequencies $\nu$ and their respective field amplitudes, we compute the pair production probability for each case. The total probability is obtained by summing the probabilities across different frequencies and then squaring the sum to derive the final result. The resulting spectrum features distinct peaks at various transverse momenta $p_{\perp}$. The peaks appear at non-zero transverse momenta, with their height and position dependent on the specific parameters of the electric field and the timing of the spectrum evaluation (see appendix A.4). The Fourier decomposition approach not only elucidates the origin of multiple peaks in the momentum spectrum but also establishes a connection between these features and the multi-photon process.
@@
Now, we can establish the specific time scales associated with the substructure observed in the TMS of created quasi-particles. By analyzing the peak locations in the transverse distribution function $f(p_{\perp},t)$, we identify three distinct time scales. Initially, after the time interval $T_{1}\approx 0.28\frac{\tau^{0.1}}{E_{0}^{2}}$, the smooth spectrum begins to reveal an interference pattern, manifesting as a multi-peak structure centered around a non-zero momentum. As time progresses, the spectrum undergoes continuous changes, characterized by oscillatory behavior around $p_{\perp}=0$. This is seen as dips and valleys splitting into multiple peaks. These oscillations persist only up to the time scale $T_{2}\approx 0.11\frac{\tau^{0.7}}{E_{0}^{5/2}}$, indicating a multiphoton signature of particle creation at finite times. After $T_{2}$, the spectrum simplifies, with a dominant peak appearing at $p_{\perp}=0$, accompanied by smaller side peaks. The maximum peak remains at $p_{\perp}=0$, while the small peaks at finite momenta gradually diminish over the time scale $T_{3}\approx 0.66\frac{\tau^{0.5}}{E_{0}^{9/2}}$. At this time, the system reaches the REPP stage, where the quasi-particle behaves like a real particle.
Next, we investigate the impact of longitudinal momentum on TMS at finite time. We observe the interesting feature that different longitudinal values manifest distinct substructures in the spectrum at a given time, as explicitly verified from Figs. 18(a)-(b). The most common feature observed is that the magnitude of the distribution function decreases for higher longitudinal momentum. Before particles reach the on-shell condition, the spectra of created particles at times $t>3\tau$ exhibit a dramatic change in structure, contingent on the value of $p_{\parallel}.$ For $p_{\parallel}=0.1[m]$ and $p_{\parallel}=0.3[m]$, the maximum peak value occurs at vanishing transverse momentum. In contrast, for $p_{\parallel}=0.5[m]$, the maximum peak is observed at non-vanishing transverse momentum, as shown in Fig. 18(b). At this finite time, we observe the sharing of longitudinal and transverse motion. The qualitative change in the shape of the spectrum, due to the influence of longitudinal momentum, can be understood mathematically. Specifically, it requires more time for the quasi-particle to evolve into a real particle with the inclusion of longitudinal momentum when analyzing the TMS of quasi-particles at finite time. In this sense, we can say that the motion of the longitudinal and transverse directions is strongly coupled, and its impact is easily seen in Figs.18. These concerns can be understood by examining the highly oscillating phase kernel $\cos{\Theta_{\bm{p}}(t)}$ present in the quantum kinetic equation (QKE) Banerjee and Singh (2019). This kernel depends on the quasi-energy $\omega(p_{\parallel},p_{\perp},t)$, which connects the longitudinal and transverse motions of the created particles. At the REPP stage, the substructure in the spectrum disappears, and the peak height for higher longitudinal momentum value is consistently lower, shown in Fig. 18(c).

### v5

@@ IV.3 Vacuum Polarisation and depolarisation function
For a better understanding of the phenomenon of particle creation under a strong electric field, we will also trace the evolution of the vacuum polarization function $u(p_{\parallel},t),$ and its counterterm $v(p_{\parallel},t)$, which governs the depolarization. Figure 13 shows the time evolution of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ for different values of $p_{\parallel}$ at zero transverse momentum. Due to pair annihilation being stronger than pair creation in the early times, the depolarisation function $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$ (See Figs. 13(a) and (c)). It appears that in the polarization function, there is a sinusoidal function behavior. On the other hand, the depolarization function $v(p_{\parallel},t)$ shows an unimodal Gaussian peak structure in its temporal evolution. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ show oscillations with varying amplitudes during the initial transient stage. These oscillations are particularly pronounced when $p_{\parallel}=0[m]$, as evident from Figs.13(a) and (c). As time progresses, irregular oscillations are observed in the transient stage. However, the oscillations become regular and stable as the system enters the REPP stage. Both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ now exhibit regular oscillations centered around the zero value. Moreover, as the momentum value $p_{\parallel}$ increases, the amplitudes of these oscillations diminish, as shown in Fig. 13(b) and (d). During the REPP stage, one interesting finding is that $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ demonstrate balancing behavior characterized by similar oscillatory patterns. This balance results from the formation of real electron-positron pairs.
@@ IV.3.1 Momentum Spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$
The longitudinal momentum significantly impacts the vacuum polarization function’s qualitative traits. To understand its dependence on $p_{\parallel}$, we comprehensively analyze the momentum spectra of both $u(p_{\parallel},t)$ and its associated counter term $v(p_{\parallel},t)$. Figure 14 shows the momentum spectra of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$ In the initial stages of particle formation, specifically at $t<0$, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ display asymmetric Gaussian peaks with roughly similar profiles. During this time, $v(p_{\parallel},t)$ dominates over $u(p_{\parallel},t)$. Upon closer examination, it becomes evident that the peaks occur at different positions around $p_{\parallel}\approx 1$, as illustrated in Fig. 14(a). At $t=0$, the initial Gaussian-like structure of $u(p_{\parallel},t)$ is deformed, becoming bi-modal asymmetric structure. The spectrum $u(p_{\parallel},t)$ exhibits a peak at $p_{\parallel}\approx-0.35\ [m]$ and a dip at $p_{\parallel}\approx 0.35\ [m]$. This indicates a significant asymmetry in the function. On the other hand, the $v(p_{\parallel},t)$ spectrum maintains its Gaussian-like unimodal profile, with a peak at $p_{\parallel}=0$. During this stage, while the shape of $u(p_{\parallel},t)$ undergoes considerable changes and becomes asymmetric, the overall form of $v(p_{\parallel},t)$ remains unchanged. Due to the force factor ``$eE(t)"$ and corresponding longitudinal quasi-momentum $P(p_{\parallel},t)=(p_{\parallel}-eA(t))$, the spectrum moved to the left side of the origin, and its peak now located at $p_{\parallel}\approx-2[m]$ for $v(p_{\parallel},t)$ whereas small dip for $u(p_{\parallel},t)$ with some disruption in the tail ( $-1<p_{\parallel}<1$) as shown in the Figs.14(c-d) near to the transient stage.
@@
At $t=24\ [m^{-1}]$, we observe two distinct structures in the polarization and depolarization functions. The first structure appears at $p_{\parallel}=-2\ [m]$ due to earlier particle creation and annihilation events. The second structure, within $-1<p_{\parallel}<1$, exhibits varying amplitude oscillations, with the maximum occurring at zero longitudinal momentum. At this stage, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ compete, as shown in Fig.14(d) and (e).
As we approach the REPP stage, $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ exhibit a single oscillating structure resembling a sine or cosine function with varying amplitude. The maximum now appears near $p_{\parallel}=0$, dominating compared to the secondary peak (and dip for $u(p_{\parallel},t)$) previously present at $p_{\parallel}\approx-2\ [m]$. As time progresses further, this left-side structure slowly vanishes, as shown in Figs. 14(e) and (f). Eventually, as $A(t)$ reaches a constant value, a balance is achieved between the processes of particle creation and annihilation, as depicted in Fig.14(g).
In the late REPP stage, where the particle distribution function $f(p_{\parallel},t)$ is constant, the polarization function $u(p_{\parallel},t)$ is balanced by its counterpart $v(p_{\parallel},t)$. This results in very regular and rapid oscillations with varying amplitude within the Gaussian envelope, as explicitly shown in Figs. 14(h-i). The overall behavior observed in both $u(p_{\parallel},t)$ and $v(p_{\parallel},t)$ during this process provides valuable insights into the complex dynamics of particle-antiparticle pair creation and annihilation under the influence of the electric field.
IV.3.2 Examining quantum interference effect in LMS: Role of Vacuum Polarization and Depolarization Functions
During the formation of electron-positron pairs from the quantum vacuum, several processes can take place. Electrons are predominantly created and annihilated simultaneously, resulting in electron acceleration and deceleration throughout the formation of real $e^{-}e^{+}$ independent pairs from the virtual $e^{-}e^{+}$ pairs. In these scenarios, the polarization functions $u(p_{\parallel},t)$ and its counterpart $v(p_{\parallel},t)$ are responsible for the acceleration and deceleration of electrons, respectively. In this context, an electron is typically created in the direction of the external field with positive momentum and is subsequently decelerated. As soon as $p_{\parallel}<0$, the electron may be annihilated again. These processes can be understood in terms of $u(p_{\parallel},t)$ and $v(p_{\parallel},t).$
@@ IV.4 Dependence on Transverse momentum
In this subsection, we extensively investigate the effect of transverse momentum on the LMS. By plotting the time evolution of the LMS for different fixed values of $p_{\perp}$, the result is displayed in Figure 16. The spectra exhibit a smooth unimodal structure during the initial stages, as depicted in Fig. 16(a). Notably, the peak’s location shifts, as illustrated in Fig.16(b), accordingly with the $(p_{\parallel}-eA(t))=0$ as explained in section IV.2. Within a narrow spectral region for $-1<p_{\parallel}<1,$ some irregular oscillating structure is observed.
Furthermore, these figures indicate that the peak value of the spectra decreases with the increase in the value of transverse momentum. Fig. 16 (c), at $t=30[m^{-1}]$ spectra shows bimodal distribution. The central peak structure with onset oscillation located at $p_{\parallel}=0$ dominates over another peak at $p_{\parallel}=-2[m]$ for a small value of $p_{\perp}.$
One interesting fact observed was that the central peak magnitude diminished very rapidly compared to other peaks, and that oscillating structure does not exist for high transverse momentum at this time. The existence of oscillation in the central peak depends on the time taken for particles to reach the on-shell configuration from the initial off-shell state, a process governed by both transverse and longitudinal momenta (see Sec.IV.1).
@@
[…]
From the occurrence of the oscillation structure, which we previously identified in the time scales discussed in section IV.2, we now observe that the quantum signature also depends on the transverse momentum, as shown in Table 3. Lower values of $p_{\perp}$ lead to oscillation patterns becoming visible earlier than those at higher values.
@@ IV.5 Transverse Momentum Spectrum
In the preceding section, we examined how fixing different values of the transverse momentum $p_{\perp}$ influences the longitudinal momentum distribution function. We now shift our attention to the transverse momentum distribution itself, focusing on the case where the longitudinal momentum is set to zero, i.e., $p_{\parallel}=0$. The transverse momentum distribution function of the created pairs at any time $t$ is denoted by $f(p_{\perp},t)$.
@@
Figure 17: TMS of created particles in the presence of time-dependent Sauter pulse at different times. The longitudinal momentum is set to zero, and all units are in electron mass units.The field parameters are $E_{0}=0.2E_{c}$ and $\tau=10[m^{-1}].$
Figure 17 shows the time evolution of the TMS for created particles at different times. At the initial time $t=-5\,[m^{-1}]$, TMS exhibits a smooth, unimodal structure with a peak at $p_{\perp}=0$, where the electric field is at $75\%$ of its maximum strength. As the electric field reaches its maximum value at $t=0$, the spectrum undergoes a dramatic change: both the peak height and width of the spectrum significantly shift, although the peak remains at $p_{\perp}=0$. After the peak reaches its maximum height, it rapidly decreases as the electric field weakens. The single peak structure initially observed at $p_{\perp}=0$ has disappeared, and new peaks have emerged at $p_{\perp}\approx\pm 0.46\,[m]$ in TMS. As an example, at $t=7\,[m^{-1}]$, TMS exhibits a bimodal profile, as shown in Fig. 17(c). As time progresses, the peaks at $p_{\perp}\approx\pm 0.46\,[m]$ tend to merge, leading to the formation of a peak at $p_{\perp}=0$ with two weakly pronounced peaks, as illustrated in Fig. 17(d). Those weakly pronounced peaks grow over time, while the height of the central peak decreases. As a result, TMS again exhibits a multi-modal structure, and $f(p_{\perp},t)$ reaches its maximum value at non-zero transverse momentum, specifically at $p_{\perp}\approx\pm 0.65\,[m]$, with a small peak still present at $p_{\perp}=0$, as shown in Fig. 17(e).
Overall, TMS displays a substructure (multiple peaks ) profile where the dominant peak is governed by non-zero values of $p_{\perp}$. Upon closer inspection, TMS reveals that the peak at $p_{\perp}=0$ becomes invisible again, and the locations of the dominant side peaks undergo slight modulation at $t=14\,[m^{-1}]$. Specifically, the dominant peak at $p_{\perp}=0.67[m]$ splits into two weakly pronounced peaks: one at $p_{\perp}=0.45[m]$ and the other at $p_{\perp}=0.94[m]$, as depicted in Fig. 17(f). As time progresses, the maximum peak at zero transverse momentum reappears. Although this central peak stands out, there are also side peaks (not shown in the figure). These side peaks eventually dominate as the central peak fades, just like in the earlier stages. In this way, the side peaks and the central peak reach the same height, as shown in the figure at $t=16.6[m^{-1}]$, and the spectrum shows a three-peak pattern.
We can say that the spectrum of created particles at a given time exhibits key features, indicating that the spectrum undergoes continuous changes and reveals two distinct structures. One scenario shows a maximum peak at $p_{\perp}=0$, suggesting particles with no transverse momentum have a high probability of creation. The other case displays a maximum peak of the spectrum at a non-zero $p_{\perp}$ value. In both cases, there are multiple peaks, showing the complex behavior of the spectrum over time, as seen in figure 17(a)-(h).
@@
The effect of created pairs tends to have a peak in the momentum distribution function at non-zero transverse momentum, as discussed by Krajewska et al. Krajewska and Kamiński (2019). It is demonstrated that the transverse momentum distribution of real pairs exhibits an off-axis maximum above the one-photon threshold frequency (or high frequency); otherwise, the maximum for the transverse distribution function occurs at $p_{\perp}=0[m]$ in the presence of an oscillating time-dependent electric field. This study gives the clue that our case of off-axis maxima or multi-modal structure may also be observed due to multiphoton transition processes. To confirm this, we analyze the probability of pair production by calculating the amplitude in the two-level equivalent formalism Krajewska and Kamiński (2019). We decompose the time-dependent electric field $E(t)$ via the Fourier transformation into cosine and sine components. This decomposition isolates individual frequency components of the field, allowing us to study their effects on pair production at finite times. For a given frequency $\nu$, the corresponding electric field amplitude $\tilde{E}(\nu)$ facilitates the evaluation of the particle production spectrum at a specific finite time $t$ (see section A.2). The spectrum exhibits multiple peaks, forming a multimodal profile. Each peak corresponds to distinct energy or momentum exchanges between the field and the particle pairs, characteristic of multi-photon absorption. By considering different frequencies $\nu$ and their respective field amplitudes, we compute the pair production probability for each case. The total probability is obtained by summing the probabilities across different frequencies and then squaring the sum to derive the final result. The resulting spectrum features distinct peaks at various transverse momenta $p_{\perp}$. The peaks appear at non-zero transverse momenta, with their height and position dependent on the specific parameters of the electric field and the timing of the spectrum evaluation (see appendix A.2). The Fourier decomposition approach not only elucidates the origin of multiple peaks in the momentum spectrum but also establishes a connection between these features and the multi-photon process.
@@
Next, we investigate the impact of longitudinal momentum on TMS at finite time. We observe the interesting feature that different longitudinal values manifest distinct substructures in the spectrum at a given time, as explicitly verified from Figs. 18(a)-(b). The most common feature observed is that the magnitude of the distribution function decreases for higher longitudinal momentum. Before particles reach the on-shell condition, the spectra of created particles at times $t>3\tau$ exhibit a dramatic change in structure, contingent on the value of $p_{\parallel}.$ For $p_{\parallel}=0.1[m]$ and $p_{\parallel}=0.3[m]$, the maximum peak value occurs at vanishing transverse momentum. In contrast, for $p_{\parallel}=0.5[m]$, the TMS peak shifts to $p_{\perp}\approx 0.6[m]$ (Fig. 18(b)), indicating energy sharing between modes. At this finite time, we observe the sharing of longitudinal and transverse motion. The qualitative change in the shape of the spectrum, due to the influence of longitudinal momentum. Specifically, it requires more time for the quasi-particle to evolve into a real particle with the inclusion of longitudinal momentum when analyzing the TMS of quasi-particles at finite time. In this sense, we can say that the motion of the longitudinal and transverse directions is strongly coupled, and its impact is easily seen in Figs.18. These concerns can be understood by examining the highly oscillating phase kernel $\cos{(\Theta_{\bm{p}}(t_{0},t))}$ present in the quantum kinetic equation (QKE) Banerjee and Singh (2019). This kernel depends on the quasi-energy $\omega(\bm{p},t)$, which connects the longitudinal and transverse motions of the created particles. At the REPP stage, the substructure in the spectrum disappears, and the peak height for higher longitudinal momentum value is consistently lower, shown in Fig. 18(c).